# What Variables Affect Out-of-Distribution Generalization in Pretrained Models?

### v2

@@ Abstract
Embeddings produced by pre-trained deep neural networks (DNNs) are widely used; however, their efficacy for downstream tasks can vary widely. We study the factors influencing out-of-distribution (OOD) generalization of pre-trained DNN embeddings through the lens of the tunnel effect hypothesis, which suggests deeper DNN layers compress representations and hinder OOD performance. Contrary to earlier work, we find the tunnel effect is not universal. Based on 10,584 linear probes, we study the conditions that mitigate the tunnel effect by varying DNN architecture, training dataset, image resolution, and augmentations. We quantify each variable’s impact using a novel SHAP analysis. Our results emphasize the danger of generalizing findings from toy datasets to broader contexts.
@@ 1 Introduction
The Tunnel Effect Hypothesis: An overparameterized $N$-layer DNN forms two distinct groups: 1. The extractor consists of the first $K$ layers, creating linearly separable representations. 2. The tunnel comprises the remaining $N-K$ layers, compressing representations and hindering OOD performance.
To test the tunnel effect hypothesis, they trained models on an in-distribution (ID) dataset and compared linear probes trained and evaluated on either ID or OOD datasets for embeddings produced by each DNN layer. They showed that ID accuracy increased monotonically, whereas OOD accuracy rapidly decreased after the extractor (Fig. 1). The likely cause of the tunnel effect is intermediate neural collapse [10]. Both [10] and [9] showed that collapsed/tunnel layers could be static without needing learning. If these results are universal, they suggest universal visual features learned in early layers and using embeddings from the penultimate layers of pre-trained DNNs, need to be rethought. However, both [9] and [10] limited their experiments to datasets with low-resolution images and relatively few categories (CIFAR-10, MNIST). Given the widespread use of embeddings from pre-trained DNNs for downstream OOD tasks, we aim to assess the universality of the tunnel effect and the variables that influence its strength.
@@
    We define metrics to measure the strength of the tunnel effect and use a SHAP-based analysis to assess each variable’s impact, e.g., image resolution, number of semantic classes, and DNN architecture. Using 64 pre-trained ID backbones and 8,604 linear probes, we identify conditions that exacerbate, mitigate, and eliminate the tunnel effect.
@@ 3.4 Statistical Analysis
To jointly analyze and rank the contribution of each variable, we use SHAP  [^1] https://github.com/shap/shap, which determines the contribution of each input variable to its output [69]. Following [70], we train Gradient Boosting Regression models to predict three output targets: a) % OOD performance retained, b) Pearson correlation, and c) ID/OOD alignment, from 8 input variables: 1) resolution, 2) augmentation, 3) ID class count, 4) spatial reduction, 5) stem, 6) CNN vs. ViT, 7) over-parametrization level, and 8) depth. We then obtain SHAP values for each variable, where using Gradient Boosting Regression facilitates controlling for variable interaction effects [70]. Because SHAP magnitude does not indicate the direction of a variable’s impact, for each of the 3 models, we fit a linear regression model between each variable and its corresponding SHAP values to obtain its slope. Positive slopes indicate the variable improves the metric. We call this SHAP Slope. Details are given in Appendix A.9.
@@ 4.1.1 Overall Findings & SHAP Analysis
Using the average % OOD performance retained across the 8 OOD datasets to analyze all 64 of our DNNs, 4 had negligible (non-existent) tunnels, 8 had weak tunnels, 13 had medium tunnels, and 39 had strong tunnels. We use intervals of [100%, 95%] for negligible, [90%, 95%) for weak, [80%, 90%) for medium, and [0%, 80%) for strong tunnel. This demonstrates that the tunnel effect is not universal and depends on variables. We next dive into the factors that influence tunnel effect strength.
@@ 4.1.3 Resolution Results
A t-SNE analysis of embeddings produced by various layers for VGGm-11 models trained on $32\times 32$ and $224\times 224$ images is given in Fig. 18, which shows that low-resolution images exhibit intermediate neural collapse [10] and representation compression to a much greater extent than high-resolution images, which is likely why many OOD detection algorithms that work well for CIFAR fail for higher-resolution ones [11]. These results highlight the dangers of extrapolating findings from low-resolution datasets to all of deep learning.
@@ 4.1.4 DNN Architecture Results
(a) # Classes (no aug)
(b) # Classes (aug)
(c) # Samples (no aug)
(d) # Samples (aug)
@@
Our SHAP analysis shows that ID class count positively impacts OOD generalization. To further analyze this, we trained VGGm-11 on different subsets of ImageNet-100 with $32\times 32$ images where we kept the training dataset fixed at 10,000 samples but varied the class counts: 10 (1000 samples per class), 50 (200 samples per class), and 100 (100 samples per class). Experiments were done with and without augmentations. As shown in Fig. 5, increasing the number of classes greatly reduces the tunnel effect (Figs. 5a and 5b). To examine the role of the dataset size, we vary the number of samples per class from 100, 200, 500, and 700 while keeping the class count constant at 100, which has a relatively small impact on the tunnel effect (Figs. 5c and 5d).
@@ 4.3 Continual Learning Results
Table 2: Continual Learning Results. The tunnel is task-specific and impacts forgetting. $E_{t}$ and $T_{t}$ denotes the extractor and tunnel, respectively, for tasks $t\in\{1,2\}$. Orange indicates original task accuracy. Red and blue indicate maximally degraded or enhanced accuracy, respectively. Fine-tuning (FT) aims to recover 1st task performance (gray); therefore, the 2nd task columns are empty.
@@
Catastrophic forgetting is a central focus in continual learning  [17, 76]. Many methods assume forgetting occurs mostly in later layers with earlier layers serving as universal features [16, 77, 78]. The original tunnel effect paper challenged this [9]. They trained VGG-19 on two tasks sequentially, where the first had half of the CIFAR-10 classes and the second the other half. After training on each task, the tunnel and extractor were saved. They found no forgetting when tunnels were swapped, indicating no learning occurred in the tunnel for task 2, and they found reduced forgetting when fine-tuning the extractor alone. These findings are worth assessing beyond CIFAR-10. We ask: What is the role of the tunnel in mitigating forgetting?
We replicated their experimental design by training ResNet-18 on ImageNet-100, where the first task had the first 50 classes and the second had the rest. After learning each task, we saved the extractor $E_{t}$, tunnel $T_{t}$, and the classification head. Tunnels $T_{1}$ and $T_{2}$ correspond to layers 14-17 ($32\times 32$) and 15-17 ($224\times 224$). We conducted this experiment with $32\times 32$ and $224\times 224$ images. We used the tunnel definition from [9] for a fair comparison. See Appendix A.7 for additional details.
@@ 5 Discussion & Conclusions
The Tunnel Effect Hypothesis: An overparameterized $N$-layer DNN forms two distinct groups: 1. The extractor consists of the first $K$ layers, creating linearly separable representations. 2. The tunnel comprises the remaining $N-K$ layers, compressing representations and hindering OOD performance. $K$ is proportional to the diversity of training inputs, where if diversity is sufficiently high, $N=K$.
Earlier works on the tunnel effect [9] and intermediate neural collapse [10] exclusively used $32\times 32$ images ID training. We found that while DNNs trained on CIFAR always exhibited the tunnel effect, the tunnel effect was greatly reduced for ImageNet-100 at higher resolutions. This discrepancy helps explain why methods validated on CIFAR and similar datasets may not generalize in many scenarios [11, 12, 13, 14, 16, 15, 17, 18]. We urge the community to use higher-resolution datasets with 100 or more categories to improve the generality of findings.
Limitations & Future Work. We rigorously studied OOD generalization on vision datasets with supervised learning. Future work should expand this analysis to non-vision and multi-modal datasets, where the tunnel effect has not yet been studied. While we studied pre-trained SSL backbones, we could not do our SHAP analysis without having more SSL backbones and potentially using more sensitive metrics, as ours had little variability for the pre-trained backbones since most did not have tunnels. That said, valuable insights for SSL could be obtained by conducting carefully controlled paired experiments, as done in our main experiments. This would require probing at least four different SSL algorithms for each variable analyzed, where training each SSL DNN would require over 10$\times$ more time than our supervised DNNs. Additionally, SSL methods employ more advanced augmentation policies than our main results, where we used random-resized crops and horizontal flips. Replicating our SHAP analysis with multiple augmentation policies could reveal whether the OOD generalization capabilities of SSL algorithms are due to their augmentation policies versus their objective functions. Lastly, identifying regularizers or other techniques that mitigate tunnel formation should be sought for continual learning methods that start from scratch using small initial sets.
@@
This work was partly supported by NSF award #2326491. The views and conclusions contained herein are those of the authors and should not be interpreted as representing any sponsor’s official policies or endorsements. We thank Junyu Chen for feedback on an early draft of this manuscript.
@@
  • Appendix D includes the list of 100 classes in the imageNet-100 dataset and confirms there is no overlap between ID and OOD datasets.
@@ Appendix A Implementation Details & Computational Resources
In this section, we use several acronyms such as WD : Weight Decay, LR : Learning Rate, GAP : Global Average Pooling, SSL : Self-Supervised Learning, SL : Supervised Learning, ID : In-Distribution, and OOD : Out-Of-Distribution. We implemented our codes in Python using PyTorch.
@@ A.8 Evaluation Metrics
Wilcoxon Signed-Rank Test. To measure the significance of the difference between two groups, we performed Wilcoxon signed-rank test between the two groups. We compared % OOD performance retained, Pearson correlation, and ID/OOD alignment metrics between each pair. We used the wilcoxon function from the spicy.stats python library.
Cliff’s Delta. The effect size (Cliff’s Delta) quantifies the statistical difference between two sets. A bigger effect size denotes a bigger difference and the order is negligible < small < medium < large. We computed Cliff’s Delta for % OOD performance retained, Pearson correlation, and ID/OOD alignment metrics to measure the effect size between the two groups. We used the cliffs_delta function from the cliffs_delta python library (https://github.com/neilernst/cliffsDelta).
@@ Appendix B Datasets
CIFAR-100 [60] dataset is similar to CIFAR-10 but with 100 classes. And, each class has 600 images. It comprises a total of 60,000 images in which training and test sets contain 50,000 and 10,000 images respectively. Note that the classes in CIFAR-100 are mutually exclusive with those in CIFAR-10. CIFAR-100 is a public dataset but does not specify any particular license (https://www.cs.toronto.edu/~kriz/cifar.html).
@@
ImageNet-100 [58] is a subset of ImageNet-1K and contains 100 ImageNet classes. It consists of 126689 training images ($224\times 224$) and 5000 test images. The dataset is available to use for research purposes under a BSD 2-Clause License: https://github.com/HobbitLong/CMC/blob/master/LICENSE. The class names of ImageNet-100 dataset are given in Sec. D.
@@
ImageNet-R incorporates distribution shifts using different artistic renditions of object classes from the original ImageNet dataset [62]. We use a variant of ImageNet-R dataset from [86]. ImageNet-R is a challenging benchmark for continual learning, transfer learning, and OOD detection. It consists of classes with different styles and intra-class diversity and thereby poses significant distribution shifts for ImageNet-1K pre-trained models [86]. It contains 200 classes, 24000 training images and 6000 test images. The dataset is available to use for research purposes under an MIT license : https://github.com/hendrycks/imagenet-r/blob/master/LICENSE.
@@ C.2.4 ID/OOD Alignment
In Table  7, we compare ID/OOD Alignment among models trained on the ImageNet-100 ID dataset with varied resolutions. More performant models obtain higher ID/OOD alignment scores across various experiments compared to less performant models.
@@ C.7 Augmentation
In this section, we study how augmentation affects the tunnel effect. We used all 512 experiments to assess the impact of augmentation. 256 experiments used augmentations and another 256 experiments did not use any augmentation. Fig. 12 exhibits that increasing augmentation greatly improves % OOD performance retained.
@@ C.8 ID Class Count
Figure 14: Class vs Data. The figure shows the trend that models trained with more classes and more samples have better OOD accuracy (%). We use C and S to denote the number of classes and number of samples per class respectively, in the X-axis of the figure.

### v3

@@ Abstract
Embeddings produced by pre-trained deep neural networks (DNNs) are widely used; however, their efficacy for downstream tasks can vary widely. We study the factors influencing transferability and out-of-distribution (OOD) generalization of pre-trained DNN embeddings through the lens of the tunnel effect hypothesis, which is closely related to intermediate neural collapse. This hypothesis suggests that deeper DNN layers compress representations and hinder OOD generalization. Contrary to earlier work, our experiments show this is not a universal phenomenon. We comprehensively investigate the impact of DNN architecture, training data, image resolution, and augmentations on transferability. We identify that training with high-resolution datasets containing many classes greatly reduces representation compression and improves transferability. Our results emphasize the danger of generalizing findings from toy datasets to broader contexts.
@@ 1 Introduction
The Tunnel Effect Hypothesis: An overparameterized $N$-layer DNN forms two distinct groups: 1. The extractor consists of the first $K$ layers, creating linearly separable representations. 2. The tunnel comprises the remaining $N-K$ layers, compressing representations and hindering OOD generalization.
To test the tunnel effect hypothesis, they trained models on an in-distribution (ID) dataset and compared linear probes trained and evaluated on either ID or OOD datasets for embeddings produced by each DNN layer. They showed that ID accuracy increased monotonically, whereas OOD accuracy rapidly decreased after the extractor (Fig. 1). The likely cause of the tunnel effect is intermediate neural collapse [10]. Both [10] and [9] showed that collapsed/tunnel layers could be static without needing learning. If these results are universal, they suggest universal visual features learned in early layers and using embeddings from the penultimate layers of pre-trained DNNs, need to be rethought. However, both [9] and [10] limited their experiments to datasets with low-resolution images and relatively few categories (CIFAR-10, MNIST). Given the widespread use of embeddings from pre-trained DNNs for downstream OOD tasks, we aim to assess the universality of the tunnel effect and the variables that influence its strengthProject website and code: https://yousuf907.github.io/oodg.
@@
    We define metrics to measure the strength of the tunnel effect and use a SHAP-based analysis to assess each variable’s impact, e.g., image resolution, number of semantic classes, and DNN architecture. Using 64 pre-trained ID backbones and 8,604 linear probes, we identify conditions that exacerbate, reduce, and eliminate the tunnel effect.
@@ 3.4 Statistical Analysis
To jointly analyze and rank the contribution of each variable, we use SHAP https://github.com/shap/shap, which determines the contribution of each input variable to its output [69, 70]. Following [71], we train Gradient Boosting Regression models to predict three output targets: a) % OOD performance retained, b) Pearson correlation, and c) ID/OOD alignment, from 8 input variables: 1) resolution, 2) augmentation, 3) ID class count, 4) spatial reduction, 5) stem, 6) CNN vs. ViT, 7) over-parametrization level, and 8) depth. We then obtain SHAP values for each variable, where using Gradient Boosting Regression facilitates controlling for variable interaction effects [71]. Because SHAP magnitude does not indicate the direction of a variable’s impact, for each of the 3 models, we fit a linear regression model between each variable and its corresponding SHAP values to obtain its slope. Positive slopes indicate the variable improves the metric. We call this SHAP Slope. Details are given in Appendix A.9.
@@ 4.1.1 Overall Findings & SHAP Analysis
Using the average % OOD performance retained across the 8 OOD datasets to analyze all 64 of our DNNs, 4 had negligible (non-existent) tunnels, 8 had weak tunnels, 13 had medium tunnels, and 39 had strong tunnels. We use intervals of [100%, 95%] for negligible, [90%, 95%) for weak, [80%, 90%) for medium, and [0%, 80%) for strong tunnel. This demonstrates that the tunnel effect is not universal and depends on variables. Next, we dive into the factors that influence tunnel effect strength.
@@ 4.1.3 Resolution Results
A t-SNE analysis for various layers from VGGm-11 models trained on $32\times 32$ and $224\times 224$ images is given in Fig. 4. The low-resolution model exhibits much greater intermediate neural collapse [10] and representation compression than the high-resolution model. This is likely why many OOD detection algorithms that work well for CIFAR fail for higher-resolution datasets [11]. These results highlight the dangers of extrapolating findings from low-resolution datasets to all of deep learning.
@@
Figure 4: High-resolution model does not exhibit representation compression. The t-SNE comparison between VGGm-11 models trained on low- (1st row) and high-resolution (2nd row) images of the same ID dataset (ImageNet-100) in an augmentation-free setting. Layer 8 marks the start of the tunnel in VGGm-11 trained on $32\times 32$ images whereas $224\times 224$ resolution does not create any tunnel. Layer 10 is the penultimate layer. The tunnel layers (layers 8-10) progressively compress representations for $32\times 32$ resolution whereas corresponding layers for $224\times 224$ resolution do not exhibit similar compression. For clarity, we show 5 classes from ImageNet-100 and indicate each class by a distinct color. The formation of distinct clusters in the $32\times 32$ model is indicative of representation compression and intermediate neural collapse [10], which impairs OOD generalization.
4.1.4 DNN Architecture Results
[…]
@@ 4.1.4 DNN Architecture Results
(a) # Classes without augmentations
(b) # Classes with augmentations
(c) # Samples without augmentations
(d) # Samples with augmentations
@@
Our SHAP analysis shows that ID class count positively impacts OOD generalization. To further analyze this, we trained VGGm-11 on different subsets of ImageNet-100 with $32\times 32$ images where we kept the training dataset fixed at 10,000 samples but varied the class counts: 10 (1000 samples per class), 50 (200 samples per class), and 100 (100 samples per class). Experiments were done with and without augmentations. As shown in Fig. 6, increasing the number of classes greatly reduces the tunnel effect (Figs. 6(a) and 6(b)). To examine the role of the dataset size, we vary the number of samples per class from 100, 200, 500, and 700 while keeping the class count constant at 100, which has a relatively small impact on the tunnel effect (Figs. 6(c) and 6(d)).
@@ 4.3 Continual Learning Results
Catastrophic forgetting is a central focus in continual learning  [17, 77]. Many methods assume forgetting occurs mostly in later layers with earlier layers serving as universal features [16, 78, 79]. The original tunnel effect paper challenged this [9]. They trained VGG-19 on two tasks sequentially, where the first task included half of the CIFAR-10 classes and the second task included the other half. After training on each task, the tunnel and extractor were saved. They found no forgetting when tunnels were swapped, indicating no learning occurred in the tunnel for task 2, and they found reduced forgetting when fine-tuning the extractor alone. These findings are worth assessing beyond CIFAR-10. We ask: What is the role of the tunnel in mitigating forgetting?
We replicated their general approach by training ResNet-18 on ImageNet-100, where the first task had the first 50 classes and the second had the rest. After learning each task $t$, we saved the extractor $E_{t}$, tunnel $T_{t}$, and the classification head. We conducted this experiment with $32\times 32$ and $224\times 224$ images. Using the tunnel definition from [9], tunnels $T_{1}$ and $T_{2}$ correspond to layers 14-17 for $32\times 32$ images and 15-17 for $224\times 224$ images. See Appendix A.7 for additional details.
Table 2: Continual Learning Results. The tunnel is task-specific and impacts forgetting. $E_{t}$ and $T_{t}$ denote the extractor and tunnel, respectively, for tasks $t\in\{1,2\}$. Orange indicates original task accuracy. Red and blue indicate degraded or maximally enhanced accuracy, respectively. Fine-tuning (FT) aims to recover 1st task performance (gray); therefore, the 2nd task columns are empty.
@@ 5 Discussion & Conclusions
The Tunnel Effect Hypothesis: An overparameterized $N$-layer DNN forms two distinct groups: 1. The extractor consists of the first $K$ layers, creating linearly separable representations. 2. The tunnel comprises the remaining $N-K$ layers, compressing representations and hindering OOD generalization. $K$ is proportional to the diversity of training inputs, where if diversity is sufficiently high, $N=K$.
Earlier works on the tunnel effect [9] and intermediate neural collapse [10] exclusively used $32\times 32$ images for ID training. We found that while DNNs trained on CIFAR always exhibited the tunnel effect, the tunnel effect was greatly reduced for ImageNet-100 at higher resolutions. This discrepancy helps explain why methods validated on CIFAR and similar datasets may not generalize in many scenarios [11, 12, 13, 14, 16, 15, 17, 18]. We urge the community to use high-resolution datasets with 100 or more categories to improve the generality of findings, especially for studies related to representation learning, neural collapse, and OOD detection/generalization.
Limitations & Future Work. While our work validates the existence of tunnels, future research should focus on developing theoretical frameworks to help explain the tunnel effect. We rigorously studied OOD generalization on vision datasets with supervised learning. Future work could study non-vision, multi-modal [80], and biased datasets [81, 82], where the tunnel effect has not yet been studied. While we studied pre-trained SSL backbones, we could not do our SHAP analysis without having more SSL backbones. Valuable insights for SSL could be obtained by conducting carefully controlled paired experiments, as done in our main experiments. This would require probing at least four different SSL algorithms for each variable analyzed, where training each SSL DNN would require over 10$\times$ more time than our supervised DNNs. Additionally, SSL methods employ more advanced augmentation policies than ours, where we used random-resized crops and horizontal flips. Replicating our SHAP analysis with multiple augmentation policies could reveal whether the OOD generalization capabilities of SSL algorithms are due to their augmentation policies versus their objective functions. Lastly, identifying regularizers or other techniques that mitigate tunnel formation should be sought for continual learning methods that start from scratch using small initial sets. This could greatly improve forward transfer, leading to more efficient continual learning methods [83, 84, 85, 86].
@@
This work was partly supported by NSF awards #2326491, #2125362, and #2317706. The views and conclusions contained herein are those of the authors and should not be interpreted as representing any sponsor’s official policies or endorsements. We thank Junyu Chen for feedback on an early draft of this manuscript.
@@
  • Appendix E reports additional statistical results for different variables and interventions.
  • Appendix F includes the list of 100 classes in the imageNet-100 dataset and confirms there is no overlap between ID and OOD datasets.
@@ Appendix A Implementation Details & Computational Resources
In this section, we use several acronyms such as WD : Weight Decay, LR : Learning Rate, GAP : Global Average Pooling, SSL : Self-Supervised Learning, SL : Supervised Learning, ID : In-Distribution, and OOD : Out-Of-Distribution. We implemented our code in Python using PyTorch.
@@ A.8 Evaluation Metrics
Wilcoxon Signed-Rank Test. To measure the significance of the difference between the two groups, we performed a Wilcoxon signed-rank test between the two groups. We compared % OOD performance retained, Pearson correlation, and ID/OOD alignment metrics between each pair. We used the wilcoxon function from the spicy.stats python library.
Cliff’s Delta. The effect size (Cliff’s Delta) quantifies the statistical difference between the two sets. A bigger effect size denotes a bigger difference and the order is negligible < small < medium < large. We computed Cliff’s Delta for % OOD performance retained, Pearson correlation, and ID/OOD alignment metrics to measure the effect size between the two groups. We used the cliffs_delta function from the cliffs_delta python library (https://github.com/neilernst/cliffsDelta).
@@ Appendix B Datasets
CIFAR-100 [60] dataset is similar to CIFAR-10 but with 100 classes. And, each class has 600 images. It comprises a total of 60,000 images. The training and test sets contain 50,000 and 10,000 images respectively. Note that the classes in CIFAR-100 are mutually exclusive with those in CIFAR-10. CIFAR-100 is a public dataset but does not specify any particular license (https://www.cs.toronto.edu/~kriz/cifar.html).
@@
ImageNet-100 [58] is a subset of ImageNet-1K and contains 100 ImageNet classes. It consists of 126689 training images ($224\times 224$) and 5000 test images. The dataset is available to use for research purposes under a BSD 2-Clause License: https://github.com/HobbitLong/CMC/blob/master/LICENSE. The class names of ImageNet-100 dataset are given in Sec. F.
@@
ImageNet-R incorporates distribution shifts using different artistic renditions of object classes from the original ImageNet dataset [62]. We use a variant of ImageNet-R dataset from [94]. ImageNet-R is a challenging benchmark for continual learning, transfer learning, and OOD detection. It consists of classes with different styles and intra-class diversity and thereby poses significant distribution shifts for ImageNet-1K pre-trained models [94]. It contains 200 classes, 24000 training images, and 6000 test images. The dataset is available to use for research purposes under an MIT license: https://github.com/hendrycks/imagenet-r/blob/master/LICENSE.
@@ C.2.4 ID/OOD Alignment
In Table  7, we compare ID/OOD Alignment among models trained on the ImageNet-100 ID dataset with varied resolutions. More performant models obtain higher ID/OOD alignment scores across various experiments compared to less performant models. The impact of different variables and interventions on the ID/OOD alignment is shown in Fig. 19 and Fig. 20.
@@ C.7 Augmentation
In this section, we study how augmentation affects the tunnel effect. We used all 512 experiments to assess the impact of augmentation. 256 experiments used augmentations and another 256 experiments did not use any augmentation. In experiments with augmentation, we used random resized crop and random horizontal flip. Fig. 13 exhibits that increasing augmentation greatly improves % OOD performance retained.
@@ C.8 ID Class Count
Figure 15: Class Count vs. Data Quantity. The figure shows the trend that models trained with more classes and more samples have better OOD accuracy (%). We use C and S to denote the number of classes and number of samples per class respectively, in the X-axis of the figure.